Intensity of light, measured in lux(lx) varies inversely as the square of the distance between the light source and the object illuminated.
A light hangs 4.8 feet above the center of a circular table 7.2 feet in diameter. If the illumination is 25 lux at the center of the table, what is the illumination at the edge of the table?

Answer by stanbon(75887)

Intensity of light, measured in lux(lx) varies inversely as the square of the distance between the light source and the object illuminated.
A light hangs 4.8 feet above the center of a circular table 7.2 feet in diameter. If the illumination is 25 lux at the center of the table, what is the illumination at the edge of the table?
I=k/d^2
25=k/4.8^2
k=576
Equation:
I=576/d^2
Find distance from light to edge of table.
Use Pytagoras: 4.8^2 + 3.6^2 = d^2
d^2=36
Intensity = 576/36= 16 lumins
Cheers,
Stan H.